# Inching toward the QCD Axions with Axion Magnetic Resonance in Helioscopes

### v1

@@ 1.1 Overview
The importance of laboratory experiments that are independent of the dark matter abundance in the search for the QCD axion cannot be overstated. Although astrophysical and cosmological studies Marsh:2015xka ; Caputo:2024oqc put efforts towards the QCD axion parameters, and axion haloscopes ADMX:2021nhd ; CAPP:2024dtx ; HAYSTAC:2023cam ; Alesini:2020vny ; Quiskamp:2022pks even reach the QCD axion lines represented by the KSVZ Kim:1979if ; Shifman:1979if and DFSZ Dine:1981rt ; Zhitnitsky:1980tq axion models for some mass windows around $m_{a}\sim\mu{\rm eV}$, those laboratory experiments not relying on the axion dark matter would provide the most definitive evidence for the QCD axion with minimal uncertainties. The QCD axions of the KSVZ and DFSZ models remain largely unexplored over the range $m_{a}\simeq 10^{-5}\,{\rm eV}-10^{0}\,{\rm eV}$. Given that, we have proposed a resonance from a magnetic field configuration with a specific length scale Seong:2023ran , named axion magnetic resonance (AMR), to extend the axion mass range towards the QCD axion lines in light-shining-through-the-wall (LSTW) experiments such as ALPS II Bahre:2013ywa ; Ortiz:2020tgs .
@@
In this work, we propose to apply the newly identified resonance, AMR, to the axion helioscopes, e.g., CAST, BabyIAXO, IAXO, and IAXO+ Kotthaus:2005zg ; Kuster:2007ue ; CAST:2017uph ; IAXO:2019mpb ; IAXO:2020wwp , by leveraging a helical magnetic field profile. We aim to overcome the axion mass suppression caused by the growing mismatch of the dispersion relations of axions and photons, which starts to appear around $m_{a}\simeq 10^{-2}\,{\rm eV}$, and to extend the coverage towards the various realizations of the QCD axion lines, for $m_{a}\simeq\mathcal{O}(10^{-1})\,{\rm eV}$. It is worth noting that the AMR in axion helioscopes manifests differently compared to that in LSTW experiments, in the context of the energy spectrum. The LSTW experiments use a mono-chromatic energy spectrum, but the axion helioscopes use a continuum energy spectrum of solar axions. Moreover, in the DFSZ axion model, the axion-electron coupling contribution to solar axions dominates over the axion-photon coupling contribution. We thus present our constraints for the QCD axion models by taking this axion-electron coupling contribution into account. In addition, other channels through the axion couplings to electrons and nucleons are thoroughly examined as well.
@@ 1.2 Lightning Review of AMR
where $\tilde{F}^{\mu\nu}=\epsilon^{\mu\nu\alpha\beta}F_{\alpha\beta}/2$ is the dual of the field strength tensor $F_{\alpha\beta}=\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha}$. Axion-photon mixing in a background magnetic field, like all two-level quantum systems that experience the Rabi oscillation, can experience a resonantly enhanced conversion to one another if the mixing term has an explicit dependence on time. [^1] Since both particles are relativistic, time and distance are not distinguished here. Ways to induce this resonance include utilizing an external magnetic field with a harmonic spatial profile in its magnitude, a spatial helical profile in its orientation Sikivie:1983ip ; Seong:2023ran ; or a temporal oscillation Sharifian:2021vsg ; Seong:2023ran . Let us take the spatial helical profile as an example and denote the magnetic field rotation frequency to be $\dot{\theta}=d\theta/dz$ with $\theta$ being the helical angle, $z$ the propagation distance. The axion-to-photon conversion probability is given by
$P_{a{\rightarrow}\gamma}$ $\simeq\sum_{i=\pm}\frac{(g_{a\gamma}B/\sqrt{2})^{2}}{\Delta_{i}^{2}}\sin^{2}\left(\frac{\Delta_{i}l}{2}\right)\,$ (2)
@@
$\Delta_{\pm}=\sqrt{\left(\Delta m_{a\gamma}^{2}/2\omega\pm\dot{\theta}\right)^{2}+\left(g_{a\gamma}B/\sqrt{2}\right)^{2}}\,,$ (3)
where we sum over the two photon helicities. In the heavy axion mass regime where the axion-photon oscillation length $\sim 4\pi\omega/m_{a}^{2}$ becomes shorter than the size of the magnetic field domain $L_{\rm B}$, we observe that the axion-photon conversion at the AMR energy of $\omega_{a}=m_{a}^{2}/(2\dot{\theta})$ leads to a considerable enhancement in the conversion probability.
@@
We get $\Delta\omega\approx(2/\dot{\theta}L_{B})\omega_{a}$. More practically, we can require the enhancement is actually an enhancement. Imagine that we choose a helical frequency $\dot{\theta}$ that poorly matches $\Delta m_{a\gamma}^{2}/2\omega$. The presence of $\dot{\theta}$ itself can impede the axion-photon conversion. By requiring the conversion probability is not worsened by the presence of a $\dot{\theta}$, we get the width during which the experimental reach is enhanced by introducing a variation of the magnetic field.
@@
This simply gives us $\delta\omega\approx\omega_{a}$. This is usually much wider than $\Delta\omega$ since the resonance usually enhances $P_{a\rightarrow\gamma}$ by a few orders of magnitude. This can also be translated to the range of axion mass an AMR setup is sensitive to. In laser-based setups (e.g., LSTW), where a monochromatic frequency around the optical wavelength range $(\sim{\rm eV})$ is considered, AMR impacts only the specific axion mass $m_{a}=\sqrt{2\omega\dot{\theta}}$ [^2] In Ref. Seong:2023ran , a possible scanning method is proposed by exploiting a discrete domain methodology, which improves the original proposal based on alternating wigglers Sikivie:1983ip .
Unlike the mono-energetic axion flux produced from high intensity lasers, solar axions have a continuum energy spectrum. This allows a single helical setup, i.e. fixed $\dot{\theta}$, to cover a wider axion mass range as long where the resonance condition $m_{a}^{2}/2\omega=\dot{\theta}$ is satisfied at different energies for axions with different masses. we argue in this work that AMR can be equally efficient with continuum axion flux such as the solar axions. In particular, a fixed $\dot{\theta}$ will be able to enhance the sensitivity of a helioscope for a range of axion masses.
@@ 3.3 Helioscope Sensitivities on QCD Axions with AMR
The left panel of Fig. 4 illustrates the helioscope sensitivities for the KSVZ model with $E/N=0$, corresponding to the yellow line. Since the axion-electron coupling is highly suppressed in the KSVZ model due to the absence of tree-level UV contributions, solar axions from electron-induced processes are negligible compared to those from the Primakoff process, and only the axion-photon coupling is relevant across the parameter space. Therefore, the KSVZ model provides almost the same constraints as those shown in Fig. 3, where only the axion-photon coupling is considered, up to only negligible differences due to the simplified assumption in Eq. (10) (see Fig. 5).666 Since the constraints on $g_{a\gamma}$ in the literature do not account for the axion-electron coupling contribution, we tolerate the negligible differences rather than multiplying the enhancement factor $\xi(m_{a})$ by the reference values from the literature, as done for AMR in Sec. 2.2.
@@ Acknowledgements.
HS is supported by the Deutsche Forschungsgemeinschaft under Germany Excellence Strategy — EXC 2121 “Quantum Universe” — 390833306. HS appreciates the hospitality of LAPTh during the visit. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented in this article was supported by the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory under projects 20220135DR and 20230047DR. SY was supported by IBS under the project code, IBS-R018-D1
@@ Appendix A Extended Maximum Likelihood
Now imagine that an experiment has observed an ensemble of events labeled as $\mathcal{X}=\{1,2,...,N_{\rm tot}\}$. Each index correspond to a unique event. Let us imagine binning the data set such that within each bin there is only one event. Within each bin the number counting follows the Poisson distribution $P(\lambda,n)$, where $\lambda$ is the expected number of events (i.e. mean) and $n$ the actual number of events. Since the measured number of events is always unity due to the aforementioned binning method, the likelihood for each bin is given by
$\mathcal{L}_{i}$ $=P(\lambda;1)=\lambda\;\mathrm{e}^{-\lambda},$ (19)
@@
$\mathcal{L}$ $=\prod_{i}^{N_{\rm tot}}\mathcal{L}_{i}$ (21)
@@
In the second step, we approximate that $\mathcal{F}_{\rm th}(E_{k})\simeq\bar{\mathcal{F}}_{\rm th}(E_{i})$. Given the data, $N_{\rm tot}$ is constant with respect to the parameters of the model, as well as $A$, $\Delta t$, $\Delta E_{\rm bin}$, and $dE$. The last constant term can be dropped even for analyzing the errors and finding the confidence level of the parameters.
@@ Appendix B Comparing with Official Constraints
However, we stress that even such a small mismatch does not propagate into our final results. This is because the small difference only shifts the overall magnitude instead of the shape of the contours. As a result, it will not affect the AMR’s enhancement factor $\xi(m_{a})$ as defined in Eq. (7). When we apply the enhancement factor to the official results to produce the results in Figs. 3 and 4, the overall shift is effectively factorized out.

### v2

@@ 1.1 Overview
The importance of laboratory experiments that are independent of the dark matter abundance in the search for the QCD axion cannot be overstated. Although astrophysical and cosmological studies Marsh:2015xka ; Caputo:2024oqc put efforts towards the QCD axion parameters, and axion haloscopes ADMX:2021nhd ; CAPP:2024dtx ; HAYSTAC:2023cam ; Alesini:2020vny ; Quiskamp:2022pks even reach the QCD axion lines represented by the KSVZ Kim:1979if ; Shifman:1979if and DFSZ Dine:1981rt ; Zhitnitsky:1980tq axion models for some mass windows around $m_{a}\sim\mu{\rm eV}$, those laboratory experiments not relying on the axion dark matter would provide the most definitive evidence for the QCD axion with minimal uncertainties. The QCD axions of the KSVZ and DFSZ models remain largely unexplored over the range $m_{a}\simeq 10^{-5}\,{\rm eV}-10^{0}\,{\rm eV}$. Given that, we have proposed a resonance from a magnetic field configuration with a specific length scale Seong:2023ran , named axion magnetic resonance (AMR). The length scale of a magnetic field configuration, e.g., the period of spatially-varying magnetic fields, acts as an effective photon mass that cancels the axion mass and leads to the resonant conversion between axions and photons. This resonance allows light-shining-through-the-wall (LSTW) experiments, such as ALPS II Bahre:2013ywa ; Ortiz:2020tgs , to extend their axion mass coverage closer to the QCD axion lines. [^1] Branching from the original conception of the LSTW experiment VanBibber:1987rq ; Anselm:1985obz ; Anselm:1985oaj , numerous variations have been proposed, including the use of an alternating magnet design VanBibber:1987rq ; Arias:2010bha ; Arias:2010bh , optical resonators that inspired the ALPS II experiment Hoogeveen:1990vq ; Sikivie:2007qm ; Mueller:2009wt , phase-shift plates Jaeckel:2007gk , superconducting radio-frequency cavities Janish:2019dpr , a collider setup Kling:2022ehv , and the ultimate configuration of the LSTW experiment Hoof:2024gfk . In a slightly different but closely related context, ideas such as measuring a phase shift with varying magnetic field profiles Zarei:2019sva ; Sharifian:2021vsg and ALP production via synchrotron radiation Yin:2024rjb have also been explored.
@@
In this work, we propose to apply the newly identified resonance, AMR, to the axion helioscopes, e.g., CAST, BabyIAXO, IAXO, and IAXO+ Kotthaus:2005zg ; Kuster:2007ue ; CAST:2017uph ; IAXO:2019mpb ; IAXO:2020wwp , by leveraging a helical magnetic field profile. [^2] A sinusoidal magnetic field profile for axion helioscopes was discussed in Sikivie’s pioneering paper Sikivie:1983ip . We reformulate this idea by focusing on a helical magnetic field profile, which maximizes axion-photon conversion at the specific axion mass where AMR occurs. We aim to overcome the axion mass suppression caused by the growing mismatch of the dispersion relations of axions and photons, which starts to appear around $m_{a}\simeq 10^{-2}\,{\rm eV}$, and to extend the coverage towards the various realizations of the QCD axion lines, for $m_{a}\simeq\mathcal{O}(10^{-1})\,{\rm eV}$. It is worth noting that the AMR in axion helioscopes manifests differently compared to that in LSTW experiments, in the context of the energy spectrum. The LSTW experiments use a mono-chromatic energy spectrum, but the axion helioscopes use a continuum energy spectrum of solar axions. Moreover, in the DFSZ axion model, the axion-electron coupling contribution to solar axions dominates over the axion-photon coupling contribution. We thus present our constraints for the QCD axion models by taking this axion-electron coupling contribution into account. In addition, other channels through the axion couplings to electrons and nucleons are thoroughly examined as well.
@@ 1.2 Lightning Review of AMR
where $\tilde{F}^{\mu\nu}=\epsilon^{\mu\nu\alpha\beta}F_{\alpha\beta}/2$ is the dual of the field strength tensor $F_{\alpha\beta}=\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha}$. For relativistic axions and photons, the equation of motion reads
$i\partial_{z}\begin{bmatrix}\gamma_{x}\\ \gamma_{y}\\ a\end{bmatrix}$ $=\frac{1}{2\omega}\begin{bmatrix}0&0&g_{a\gamma}\omega B_{x}(z)\\ 0&0&g_{a\gamma}\omega B_{y}(z)\\ g_{a\gamma}\omega B_{x}(z)&g_{a\gamma}\omega B_{y}(z)&m^{2}_{a}\end{bmatrix}\begin{bmatrix}\gamma_{x}\\ \gamma_{y}\\ a\end{bmatrix},$
where the $z$-axis is the propagating direction, and $\gamma_{i=x,y}$ and $a$ denote the photon and axion states, respectively. $\omega$ is the energy of the axion and the photon, and $m_{a}$ is the axion mass. $B_{i=x,y}(z)$ represents the $i=x,y$ components of the background magnetic field, which vary along the $z$-direction. This set of equations describes the mixing between axions and photons in the background magnetic fields. As two-level quantum systems experience the Rabi oscillation, this axion-photon system can experience a resonantly enhanced conversion to one another if the mixing term has an explicit dependence on time. [^3] Since both particles are relativistic, time and distance are not distinguished here. Ways to induce this resonance include utilizing an external magnetic field with a harmonic spatial profile in its magnitude, a spatial helical profile in its orientation Sikivie:1983ip ; Seong:2023ran ; or a temporal oscillation Sharifian:2021vsg ; Seong:2023ran . Let us take the spatial helical profile as an example and parametrize the magnetic field components $B_{x}=B\sin\dot{\theta}z$, $B_{y}=B\cos\dot{\theta}z$ with a constant rotation frequency $\dot{\theta}$ and a magnitude $B$. If $m_{a}^{2}>g_{a\gamma}\omega B$, which is indeed the case of our interest, the axion-to-photon conversion probability is given by
$P_{a{\rightarrow}\gamma}$ $\simeq\sum_{i=\pm}\frac{(g_{a\gamma}B/\sqrt{2})^{2}}{\Delta_{i}^{2}}\sin^{2}\left(\frac{\Delta_{i}z}{2}\right)\,$ (2)
@@
$\Delta_{\pm}=\sqrt{\left(m_{a}^{2}/2\omega\pm\dot{\theta}\right)^{2}+\left(g_{a\gamma}B/\sqrt{2}\right)^{2}}\,,$ (3)
where we sum over the two photon helicities. In the heavy axion mass regime where the axion-photon oscillation length $\sim 4\pi\omega/m_{a}^{2}$ becomes shorter than the size of the magnetic field domain $L_{\rm B}$, the axion-photon conversion is suppressed by $m_{a}^{-4}$, and we observe that, at the specific energy of $\omega_{a}\equiv|m_{a}^{2}/(2\dot{\theta})|$, the conversion probability is significantly enhanced due to the cancellation between $m_{a}^{2}/(2\omega_{a})$ and ±$\pm\dot{\theta}$.
@@
We get $\Delta\omega\approx(2/\dot{\theta}L_{B})\omega_{a}$. More practically, we can require the enhancement is actually an enhancement. Imagine that we choose a helical frequency $\dot{\theta}$ that poorly matches $m_{a}^{2}/2\omega$. The presence of $\dot{\theta}$ itself can impede the axion-photon conversion. By requiring the conversion probability is not worsened by the presence of a $\dot{\theta}$, we get the width during which the experimental reach is enhanced by introducing a variation of the magnetic field.
@@
This simply gives us $\delta\omega\approx\omega_{a}$. This is usually much wider than $\Delta\omega$ since the resonance usually enhances $P_{a\rightarrow\gamma}$ by a few orders of magnitude. This can also be translated to the range of axion mass an AMR setup is sensitive to. In laser-based setups (e.g., LSTW), where a monochromatic frequency around the optical wavelength range $(\sim{\rm eV})$ is considered, AMR impacts only the specific axion mass $m_{a}=\sqrt{2\omega|\dot{\theta}|}$ [^4] In Ref. Seong:2023ran , a possible scanning method is proposed by exploiting a discrete domain methodology, which improves the original proposal based on alternating wigglers Sikivie:1983ip .
Unlike the mono-energetic axion flux produced from high intensity lasers, solar axions have a continuum energy spectrum. This allows a single helical setup, i.e. fixed $\dot{\theta}$, to cover a wider axion mass range as long where the resonance condition $m_{a}^{2}/(2\omega)=|\dot{\theta}|$ is satisfied at different energies for axions with different masses. We argue in this work that AMR can be equally efficient with continuum axion flux such as the solar axions. In particular, a fixed $\dot{\theta}$ will be able to enhance the sensitivity of a helioscope for a range of axion masses.
@@ 2.2 Impact of AMR on Helioscope Sensitivity
2.3 Experimental Relevance
Figure 4: Schematic picture of helical currents (left), within which generate helical magnetic fields around the axis. The right panels represent the resulting magnetic field profiles for helical currents, depending on the parameters of helical currents. The results show that the $B_{x}$ (red) and $B_{y}$ (green) components reproduce the sine and cosine forms (solid lines) well. The $B_{z}$ component is colored blue. The relevant parameters are defined with respect to the radius $R$. (Right top) For $R\,\Omega=0.1$, the magnetic field is calculated at the radial coordinate $\rho=0.3R$. The number of current cables, $N_{\rm cab}$, is set to $5$, within a bundle width of approximately half the spatial period of the current, $0.5\times(2\pi/\Omega)\times(N_{\rm cab}-1)/N_{\rm cab}$. (Right bottom) Same as the top panel, but with $R\,\Omega=3.0$.
Since we utilize spatially varying magnetic fields, it is essential to establish the feasibility of these configurations before advancing to the next section. First of all, mature technologies for helical magnets, such as those used in the RHIC Snake magnet zotero-21781 ; zotero-21778 ; zotero-21777 ; Anerella:2003se , already exist. Likewise, one way to realize a helical magnetic field profile is employing helical currents. To demonstrate this idea, we calculate the magnetic field profiles generated by helical currents, as shown in Fig. 4, using the Biot-Savart law. The position vector for a single helical current can be parametrized as $\mathbf{l^{\prime}}=(R\cos\phi^{\prime},R\sin\phi^{\prime},\phi^{\prime}/\Omega)$, where $\Omega$ is the spatial frequency of the helical currents, resulting in the magnetic fields with $\dot{\theta}=\Omega$. We can factor out the radius $R$ and define all length scales relative to $R$. Up to normalization, we find that the magnetic fields exhibit a helical profile around the central axis of the currents. To broaden the region where the magnetic fields closely follow the helical profile, we introduce additional currents to form a helical bundle. As shown in Fig. 4, at the radial coordinate $\rho=0.3R$, we find the magnetic fields still follow the helical profile. To satisfy Maxwell’s equations, the $B_{z}=B_{z}(x,y,z)$ component appears, which does not impact our sensitivity.
We also examine a discrete setup to relax the requirements. We define a discrete helical array with an integer $N$ as a series of constant magnetic field units having rotated by a fixed angle of $2\pi/N$ relative to the previous one. In this definition, alternating dipole magnets, referred to as wigglers or undulators, correspond to a specific case of $N=2$. The photon spectrum converted from solar axions as Fig. 1 is derived for the magnetic fields of a discrete helical array with $N=3$ in Fig. 5. The results indicate that the $N=3$ configuration already produces promising signals comparable to those of the continuous helical magnetic field configuration. Even we observe that the $N=2$ case yields sufficiently strong signals since the magnetic fields contain a significant helical contribution within their Fourier modes. To sum up, both continuous and discrete realizations of helical magnetic fields exist and provide robust benchmarks to adopt these setups for our purpose.
Figure 5: Similar to Fig. 1, but showing the case of a discrete helical array with $N=3$. For this configuration, the constant magnetic field is rotated by $2\pi/N$ at every $(2\,{\rm m})/N$ along the propagation direction to mimic the continuous helical profile.
@@ 3.3 Helioscope Sensitivities on QCD Axions with AMR
The left panel of Fig. 6 illustrates the helioscope sensitivities for the KSVZ model with $E/N=0$, corresponding to the yellow line. Since the axion-electron coupling is highly suppressed in the KSVZ model due to the absence of tree-level UV contributions, solar axions from electron-induced processes are negligible compared to those from the Primakoff process, and only the axion-photon coupling is relevant across the parameter space. Therefore, the KSVZ model provides almost the same constraints as those shown in Fig. 3, where only the axion-photon coupling is considered, up to only negligible differences due to the simplified assumption in Eq. (10) (see Fig. 7).888 Since the constraints on $g_{a\gamma}$ in the literature do not account for the axion-electron coupling contribution, we tolerate the negligible differences rather than multiplying the enhancement factor $\xi(m_{a})$ by the reference values from the literature, as done for AMR in Sec. 2.2.
@@ Acknowledgements.
HS is supported by the Deutsche Forschungsgemeinschaft under Germany Excellence Strategy — EXC 2121 “Quantum Universe” — 390833306. HS appreciates the hospitality of LAPTh during the visit. This work was partly supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented in this article was supported by the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory under projects 20220135DR and 20230047DR. CS acknowledges the receipt of the grant from the Abdus Salam International Centre for Theoretical Physics (ICTP), Trieste, Italy, where part of this work was finished. CS thanks the hospitality and support of IBS-CTPU and support of CKC for the CTPU-CKC Joint Focus Program: Let there be light (particles) Workshop, where part of this work was performed. SY was supported by IBS under the project code, IBS-R018-D1
@@ Appendix A Extended Maximum Likelihood
Now imagine that an experiment has observed an ensemble of events labeled as $\mathcal{X}=\{1,2,...,N_{\rm tot}\}$. Each index corresponds to a unique event. Let us imagine binning the data set such that within each bin there is at most one event. Within each bin the number counting follows the Poisson distribution $P(\lambda,n)$, where $\lambda$ is the expected number of events (i.e. mean) and $n$ the actual number of events. Since the measured number of events is always zero or unity due to the aforementioned binning method, the likelihood for each bin is given by
$\mathcal{L}_{j}^{0}$ $=P(\lambda;0)=\mathrm{e}^{-\lambda},\qquad\mathcal{L}_{k}^{1}=P(\lambda;1)=\lambda\;\mathrm{e}^{-\lambda},$ (19)
@@
$\mathcal{L}$ $=\left(\prod_{j}\mathcal{L}_{j}^{0}\right)\left(\prod_{k}^{N_{\rm tot}}\mathcal{L}_{k}^{1}\right)$ (21)
@@
In the first equality, we approximate that $\mathcal{F}_{\rm th}(E_{k})\simeq\bar{\mathcal{F}}_{\rm th}(E_{i})$. Given the data, $N_{\rm tot}$ is constant with respect to the parameters of the model, as well as $A$, $\Delta t$, $\Delta E_{\rm bin}$, and $dE$. The last constant term can be dropped even for analyzing the errors and finding the confidence level of the parameters.
@@ Appendix B Comparing with Official Constraints
However, we stress that even such a small mismatch does not propagate into our final results. This is because the small difference only shifts the overall magnitude instead of the shape of the contours. As a result, it will not affect the AMR’s enhancement factor $\xi(m_{a})$ as defined in Eq. (7). When we apply the enhancement factor to the official results to produce the results in Figs. 3 and 6, the overall shift is effectively factorized out.